# Little By Little: Continual Learning via Self-Activated Sparse Mixture-of-Rank Adaptive Learning

### v1

@@ Appendix B Additional Details and Results of Experiments
Table 6: Details of the 15 datasets used in our continual-learning experiments using LLM. NLI denotes natural language inference, and QA denotes question-answering tasks. The first five tasks comprise the standard CL benchmark; the remaining ten tasks are used for the extended long-sequence evaluations.
@@ 15. IMDB SuperGLUE sentiment analysis movie reviews accuracy
Table 8: The six task-sequence orders used in our continual learning experiments. Sequences 1–3 follow the standard CL benchmarks employed in prior work. Sequences 4–6 extend to longer 15-task streams, as introduced in [62].
@@
B.1 Detailed Experiment Settings
@@
CL on CLIP. We follow the experimental setups in [92, 85, 81] and use the CLIP model with a ViT-B/16 backbone [60] for all experiments. By default, MoRA is applied to every pre-trained weight matrix in both the vision and text encoders, with an initial rank of 16 per update. Each task is trained for 500 iterations using AdamW [47] with learning rate of $45e-$ 4. During continual learning, we freeze the ranks learned from previous tasks and initialize new $r=16$ ranks for each incoming task. The rank activation budget is set to 16 throughout all tasks. We set the temperature $\tau_{\text{MoRA}}=0.1$ and the threshold $\delta=0.2$.
@@
CL on LLMs. We follow the protocol of previous work in continually fine-tuning a T5-large model on a suite of text-classification tasks. We train on five standard benchmarks—AG News, Amazon Reviews, Yelp Reviews, DBpedia, and Yahoo Answers—using three distinct task orderings drawn from [59, 62, 74, 58]. To probe longer sequences, we extend this to a 15-dataset stream (Table 6), incorporating tasks from the original CL benchmark [90], GLUE [68], SuperGLUE [69], and the IMDB movie reviews corpus. Natural language prompts for each task are presented in Table 7, with NLI tasks (MNLI, RTE, CB), sentiment classification (Amazon, Yelp, SST-2, IMDB), and topic classification (AG News, DBpedia, Yahoo).
@@
We evaluate three distinct task sequences for both the standard CL and 15-task benchmarks (Table 8). After completing the final task in each stream, we report the average accuracy across all tasks. All experiments use one epoch per task with DeepSpeed, a fixed learning rate of $31e-$ 3, batch size 64, and dropout of 0.1. MoRA is applied to both the query and key projection matrices within attention layers, initializing $r=8$ new ranks for each incoming task similarly as in [74, 58]. We maintain a constant activation budget of 4 ranks throughout continual learning, set the temperature $\tau_{\text{MoRA}}=0.1$, and the threshold $\delta=0.2$.
@@
Generalization and forgetting on unseen tasks. To assess effects on pre-trained general knowledge, we fine-tune Llama3.1-8B [28] on the CodeAlpaca code-generation dataset [9] using llama-Factory [91] and evaluate using lm-eval-harness [26] on zero-shot in-domain performance on HumanEval [12], as well as out-of-domain accuracy on a broad selection of MMLU [31] subjects—Formal Logic, Philosophy, World Religions, Economics, Public Relations, STEM, Physics, and Machine Learning.
@@
In this experiment, MoRA is applied to all linear weight matrices of the pre-trained model. We fine-tune on CodeAlpaca with a batch size of 32 over 3 epochs and a cosine learning-rate schedule starting at $45e-$ 4. We train with $r=16$ ranks and enforce a constant activation budget of 4 ranks. The self-routed gating uses a temperature $\tau_{\text{MoRA}}=0.5$ and a threshold $\delta=0.2$. We observe that, due to variations in hidden representations across architectures, the optimal temperature setting can differ across each pre-trained models.
Reproducibility Statement. Our source code will be publicly released upon paper acceptance. Results are averaged over three independent runs to ensure robustness. Experiments with the CLIP model were carried out on an NVIDIA GeForce RTX 3090 GPU, while experiments with LLMs utilized 4 NVIDIA H100 GPUs.
@@
Table 9: Average accuracy on T5-large continual-learning benchmarks after the final task, evaluated over extended 15-task sequences. Results for prior methods are taken from [58].
@@
B.2 Continual Learning of LLM on Long Task Sequences
In Sec. 4.2 Table 3, we evaluate MoRA on standard CL benchmarks [62]. In Table 9, we extend the evaluation to challenged long task sequences using 15 datasets with 3 different orderings as in [74, 58]. Consistent with the findings in Table 3, MoRA outperforms previous methods in terms of averaged performance across three task orders, and largely close the gap to multi-task learning. Two key design choices drive this robust performance in the long-sequence regime. First, by decomposing each LoRA update into fine-grained rank-1 components and enforcing a small, fixed activation budget, MoRA encourages each rank to specialize on a narrow subspace of the data manifold. At inference time, only the most relevant subspaces are activated for a given input, which preserves earlier task representations and prevents catastrophic interference. Second, our self-routed gating mechanism enables each rank to assess its own relevance on a per-token basis, yielding stable mixture patterns as the expert pool grows. Coupling with our proposed rank pruning, these mechanisms ensure that MoRA continually incorporates new knowledge only when needed while robustly maintaining prior capabilities.
@@ B.2 Continual Learning of LLM on Long Task Sequences
LoRA [32] $r\,(d_{\text{in}}+d_{\text{out}})$
@@
MoE-Adapter (2 experts/task) [85] $2(r\,(d_{\text{in}}+d_{\text{out}})+d_{\text{in}})$
CoDyRA [48] $r\,(d_{\text{in}}+d_{\text{out}})$
O-LoRA [74] $r\,(d_{\text{in}}+d_{\text{out}})$
LB-CL [58] $r\,(d_{\text{in}}+d_{\text{out}})+r$
@@
B.3 Computation Cost

### v2

@@ Appendix A Additional Details and Results of Experiments
Table 5: Details of the 15 datasets used in our continual‐learning experiments using LMs. NLI denotes natural language inference, and QA denotes question‐answering tasks. The first five tasks comprise the standard CL benchmark; the remaining ten tasks are used for the extended long-sequence evaluations.
@@ 15. IMDB SuperGLUE sentiment analysis movie reviews accuracy
Table 7: The six task‐sequence orders used in our continual learning experiments. Sequences 1–3 follow the standard CL benchmarks employed in prior work. Sequences 4–6 extend to longer 15‐task streams, as introduced in (Razdaibiedina et al., 2023).
@@
A.1 Detailed Experiment Settings
@@
CL on CLIP. We follow the experimental setups in (Zheng et al., 2023; Yu et al., 2024; Xu et al., 2024) and use the CLIP model with a ViT-B/16 backbone (Radford et al., 2021) for all experiments. By default, MoRA is applied to every pre-trained weight matrix in both the vision and text encoders, with an initial rank of 16 per update. Each task is trained for 500 iterations using AdamW (Loshchilov & Hutter, 2017) with a learning rate of 5e45e-4. During continual learning, we freeze the ranks learned from previous tasks and initialize new $r=16$ ranks for each incoming task. The rank activation budget is set to 16 throughout all tasks. We set the temperature $\tau_{\text{MoRA}}=0.1$ and the threshold $\delta=0.2$.
@@
CL on LMs. We follow the protocol of previous work in continually fine-tuning the T5-large (Raffel et al., 2020) and LLaMA2-7B (Touvron et al., 2023) on a suite of text-classification tasks. We train on five standard benchmarks—AG News, Amazon Reviews, Yelp Reviews, DBpedia, and Yahoo Answers—using three distinct task orderings drawn from (Qin & Joty, 2021; Razdaibiedina et al., 2023; Wang et al., 2023; Qiao & Mahdavi, 2024). To probe longer sequences, we extend this to a 15-dataset stream (Table 5), incorporating tasks from the original CL benchmark (Zhang et al., 2015), GLUE (Wang et al., 2018), SuperGLUE (Wang et al., 2019), and the IMDB movie reviews corpus. Natural language prompts for each task are presented in Table 6, with NLI tasks (MNLI, RTE, CB), sentiment classification (Amazon, Yelp, SST-2, IMDB), and topic classification (AG News, DBpedia, Yahoo).
@@
We evaluate three distinct task sequences for both the standard CL and 15-task benchmarks (Table 7). After completing the final task in each stream, we report the average accuracy across all tasks. All experiments use one epoch per task with DeepSpeed, a fixed learning rate of 1e31e-3, batch size 64, and dropout of 0.1. MoRA is applied to both the query and key projection matrices within attention layers, initializing $r=8$ new ranks for each incoming task similarly as in (Wang et al., 2023; Qiao & Mahdavi, 2024). We maintain a constant activation budget of 4 ranks throughout continual learning, set the temperature $\tau_{\text{MoRA}}=0.1$, and the threshold $\delta=0.2$.
@@
Generalization and forgetting on unseen tasks. To assess effects on pre-trained general knowledge, we fine-tune Llama3.1-8B (Grattafiori et al., 2024) on the CodeAlpaca code-generation dataset (Chaudhary, 2023) using llama-Factory (Zheng et al., 2024) and evaluate using lm-eval-harness (Gao et al., 2024) on zero-shot in-domain performance on HumanEval (Chen et al., 2021), as well as out-of-domain accuracy on a broad selection of MMLU (Hendrycks et al., 2021) subjects—Formal Logic, Philosophy, World Religions, Economics, Public Relations, STEM, Physics, and Machine Learning.
@@
In this experiment, MoRA is applied to all linear weight matrices of the pre-trained model. We fine-tune on CodeAlpaca with a batch size of 32 over 3 epochs and a cosine learning-rate schedule starting at 5e45e-4. We train with $r=16$ ranks and enforce a constant activation budget of 4 ranks. The self-routed gating uses a temperature $\tau_{\text{MoRA}}=0.5$ and a threshold $\delta=0.2$. We observe that, due to variations in hidden representations across architectures, the optimal temperature setting can differ across each pre-trained models.
Further details of the used datasets in MTIL and X-TAIL. The MTIL setting consists of 1,201 classes drawn from 11 diverse datasets: Aircraft (Maji et al., 2013), Caltech101 (Fei-Fei et al., 2004), CIFAR100 (Krizhevsky et al., 2009), DTD (Cimpoi et al., 2014), EuroSAT (Helber et al., 2019), Flowers (Nilsback & Zisserman, 2008), Food (Bossard et al., 2014), MNIST (Deng, 2012), OxfordPet (Parkhi et al., 2012), Cars (Krause et al., 2013), and SUN397 (Xiao et al., 2010). In the X-TAIL setting, a total of 10 datasets are used, with CIFAR100 (Fei-Fei et al., 2004) excluded to prevent domain overlap, following the protocol in (Xu et al., 2024). In line with (Xu et al., 2024), we use a 5-shot split for MTIL and a 16-shot split for X-TAIL.
@@
Table 8: Average accuracy on T5-large continual‐learning benchmarks after the final task, evaluated over extended 15-task sequences. Results for prior methods are taken from (Qiao & Mahdavi, 2024).
@@
Table 9: Continual learning results on standard CL benchmarks with the LLaMA2-7B model.
Method Order-1 Order-2 Order-3 Avg.
O-LoRA 76.8 75.7 75.7 76.1
MoRA 77.8 78.0 79.3 78.4
A.2 Continual Learning of LMs on Long Task Sequences
In Sec. 4.2 Table 3, we evaluate MoRA on standard CL benchmarks (Razdaibiedina et al., 2023). In Table 8, we extend the evaluation to challenged long task sequences using 15 datasets with 3 different orderings as in (Wang et al., 2023; Qiao & Mahdavi, 2024). Consistent with the findings in Table 3, MoRA outperforms previous methods in terms of averaged performance across three task orders, and largely close the gap to multi-task learning. Two key design choices drive this robust performance in the long‐sequence regime. First, by decomposing each LoRA update into fine‐grained rank-1 components and enforcing a small, fixed activation budget, MoRA encourages each rank to specialize on a narrow subspace of the data manifold. At inference time, only the most relevant subspaces are activated for a given input, which preserves earlier task representations and prevents catastrophic interference. Second, our self‐routed gating mechanism enables each rank to assess its own relevance on a per‐token basis, yielding stable mixture patterns as the expert pool grows. Coupling with our proposed rank pruning, these mechanisms ensure that MoRA continually incorporates new knowledge only when needed while robustly maintaining prior capabilities.
@@ A.2 Continual Learning of LMs on Long Task Sequences
LoRA (Hu et al., 2021) $r\,(d_{\text{in}}+d_{\text{out}})$
@@
MoE-Adapter (2 experts/task) (Yu et al., 2024) $2(r\,(d_{\text{in}}+d_{\text{out}})+d_{\text{in}})$
CoDyRA (Lu et al., 2024) $r\,(d_{\text{in}}+d_{\text{out}})+r$
O-LoRA (Wang et al., 2023) $r\,(d_{\text{in}}+d_{\text{out}})$
LB-CL (Qiao & Mahdavi, 2024) $r\,(d_{\text{in}}+d_{\text{out}})+r$
@@
A.3 Multi‐domain task incremental learning.
We evaluate MoRA in the few‐shot MTIL setting (Table 11) under the same protocols as (Yu et al., 2024; Xu et al., 2024; Lu et al., 2024). Consistent with the results observed in the X-TAIL setting, our method demonstrates clear superiority in this scenario.
In this challenging scenario, the model must learn 11 diverse tasks sequentially, with only five examples per class. These findings validate that our self-activated sparse mixture-of-ranks framework both facilitates continual acquisition of new knowledge and mitigates forgetting from the pre-trained model and earlier tasks.
Table 11: Comparisons on 5-shot MTIL setting. Following the same protocol as in (Yu et al., 2024; Xu et al., 2024; Lu et al., 2024).
Method
Aircraft Caltech101 CIFAR100 DTD EuroSAT Flowers Food MNIST OxfordPet Cars SUN397 Average
CLIP
   Zero-shot (Radford et al., 2021) 24.3 88.4 68.2 44.6 54.9 71.0 88.5 59.4 89.0 64.7 65.2 65.3
Transfer
   Zero-shot (Radford et al., 2021) 88.4 68.2 44.6 54.9 71.0 88.5 59.6 89.0 64.7 65.2 69.4
   LwF (Li & Hoiem, 2017) 72.1 49.2 35.9 44.5 41.1 66.6 50.5 69.0 19.0 51.7 50.0
   LwF-VR (Ding et al., 2022) 82.2 62.5 40.1 40.1 56.3 80.0 60.9 77.6 40.5 60.8 60.1
   WiSE-FT (Wortsman et al., 2022) 77.6 60.0 41.3 39.4 53.0 76.6 58.1 75.5 37.3 58.2 57.7
   ZSCL (Zheng et al., 2023) 84.0 68.1 44.8 46.8 63.6 84.9 61.4 81.4 55.5 62.2 65.3
   MoE-Adapter (Yu et al., 2024) 87.9 68.2 44.1 48.1 64.7 88.8 69.0 89.1 64.5 65.1 68.9
   RAIL-Primal (Xu et al., 2024) 88.4 68.2 44.6 54.9 71.0 88.5 59.6 89.0 64.7 65.2 69.4
   CoDyRA (Lu et al., 2024) 92.4 68.4 45.8 54.5 69.6 87.4 65.2 88.5 64.2 64.5 69.9
   MoRA 92.0 68.8 45.6 53.1 68.6 84.4 64.3 89.8 65.4 64.8 69.7
Average
   LwF (Li & Hoiem, 2017) 23.5 77.4 43.5 41.7 43.5 52.2 54.6 63.4 68.0 21.3 52.6 49.2
   LwF-VR (Ding et al., 2022) 24.9 89.1 64.2 53.4 54.3 70.8 79.2 66.5 79.2 44.1 61.6 62.5
   WiSE-FT (Wortsman et al., 2022) 32.0 87.7 61.0 55.8 68.1 69.3 76.8 71.5 77.6 42.0 59.3 63.7
   ZSCL (Zheng et al., 2023) 28.2 88.6 66.5 53.5 56.3 73.4 83.1 56.4 82.4 57.5 62.9 64.4
   MoE-Adapter (Yu et al., 2024) 30.0 89.6 73.9 58.7 69.3 79.3 88.1 76.5 89.1 65.3 65.8 71.4
   RAIL-Primal (Xu et al., 2024) 32.9 94.5 69.9 58.1 71.8 84.4 88.5 70.4 89.0 66.1 65.7 71.9
   CoDyRA (Lu et al., 2024) 34.6 95.8 73.9 60.0 77.1 81.3 86.6 75.9 89.9 66.1 65.3 73.3
   MoRA 36.7 95.4 74.9 61.9 77.1 82.6 85.3 76.0 90.5 67.0 65.6 73.9
Last
   LwF (Li & Hoiem, 2017) 22.1 58.2 17.9 32.1 28.1 66.7 46.0 84.3 64.1 31.5 60.1 46.5
   LwF-VR (Ding et al., 2022) 22.9 89.8 59.3 57.1 57.6 79.2 78.3 77.7 83.6 60.1 69.8 66.9
   WiSE-FT (Wortsman et al., 2022) 30.8 88.9 59.6 60.3 80.9 81.7 77.1 94.9 83.2 62.8 70.0 71.9
   ZSCL (Zheng et al., 2023) 26.8 88.5 63.7 55.7 60.2 82.1 82.6 58.6 85.9 66.7 70.4 67.4
   MoE-Adapter (Yu et al., 2024) 30.1 89.3 74.9 64.0 82.3 89.4 87.1 89.0 89.1 69.5 72.5 76.1
   RAIL-Primal (Xu et al., 2024) 32.9 95.1 70.3 63.2 81.5 95.6 88.5 89.7 89.0 72.5 71.0 77.2
   CoDyRA (Lu et al., 2024) 31.6 95.5 72.8 63.5 85.0 89.7 85.0 94.7 93.2 73.6 73.0 78.0
   MoRA 32.5 95.3 75.3 66.6 87.8 92.6 86.3 96.3 92.6 73.5 73.8 79.3
A.4 More Comparison on HumanEval for Code Generation
To further evaluate code generation performance, we compare MoRA against LoRI-D and LoRI-S (Zhang et al., 2025) on the HumanEval benchmark (Table 12). MoRA consistently outperforms both LoRI variants across all metrics, achieving notable improvements in Pass@1, Pass@5, and especially Pass@10.
Table 12: Performance comparison on the HumanEval benchmark, reported in terms of Pass@1, Pass@5, and Pass@10.
HumanEval Pass@1 Pass@5 Pass@10
LoRI-D 43.2 57.6 63.2
LoRI-S 41.3 54.4 59.6
MoRA 47.6 60.9 70.1
A.5 performance robustness
MoRA employs a sparse mixture of previously learned and newly introduced rank-1 experts to capture both shared and task-specific knowledge, resulting in substantially improved Last performance. To assess statistical significance and robustness, we report mean and standard deviation over three independent runs (Table 13). MoRA consistently outperforms competing methods across all metrics and exhibits lower variance, highlighting its effectiveness and stability in continual-learning scenarios.
Table 13: Comparison to InfLoRA and performance robustness. We report mean and standard deviation across 3 independent runs. Best performances are marked in bold.
Method
Cars Aircraft OxfordPet Food SUN397 MNIST Flowers DTD Caltech101 EuroSAT Average
Transfer
   InfLoRA $72.26^{\pm 0.56}$ $36.19^{\pm 0.64}$ $38.46^{\pm 0.39}$ $55.22^{\pm 1.65}$ $73.19^{\pm 0.55}$ $39.32^{\pm 1.54}$ $80.29^{\pm 0.91}$ $1.19^{\pm 1.16} 55$$05^{\pm 0.51} 55.$$5.69^{\pm 0.24}$
   CoDyRA $74.3^{\pm 0.52}$ $36.8^{\pm 0.23}$ $44.2^{\pm 0.56}$ $\mathbf{69.9^{\pm 0.56}}$ $\mathbf{83.5^{\pm 0.23}}$ $42.8^{\pm 0.18}$ $\mathbf{88.9^{\pm 0.42}}$ $64.6^{\pm 0.47}$ $\mathbf{63.4^{\pm 0.56}}$ $63.2^{\pm 0.28}$
   MoRA $\mathbf{74.5^{\pm 0.51}}$ $\mathbf{38.1^{\pm 0.24}}$ $\mathbf{46.9^{\pm 0.56}}$ $65.3^{\pm 0.44}$ $82.9^{\pm 0.18}$ $\mathbf{45.8^{\pm 0.31}}$ $88.2^{\pm 0.15}$ $\mathbf{65.1^{\pm 0.35}}$ $62.9^{\pm 0.10}$ $\mathbf{63.3^{\pm 0.26}}$
Average
   InfLoRA $20.49^{\pm 0.98}$ $78.58^{\pm 1.02}$ $48.5^{\pm 1.18}$ $66.59^{\pm 1.51}$ $71.83^{\pm 0.80}$ $76.79^{\pm 0.34}$ $61.45^{\pm 1.36}$ $82.59^{\pm 0.86}$ $55.3^{\pm 1.34}$ $56.67^{\pm 0.59}$ $62.48^{\pm 0.31}$
   CoDyRA $41.4^{\pm 0.28}$ $81^{\pm 0.38}$ $58.7^{\pm 0.26}$ $77.8^{\pm 0.47}$ $83.4^{\pm 0.39}$ $\mathbf{84.6^{\pm 0.28}}$ $64.5^{\pm 0.14}$ $\mathbf{90.4^{\pm 0.40}}$ $67.2^{\pm 0.23}$ $\mathbf{64.4^{\pm 0.47}}$ $71.3^{\pm 0.18}$
   MoRA $\mathbf{44.1^{\pm 0.24}}$ $\mathbf{81.6^{\pm 0.34}}$ $\mathbf{64.6^{\pm 0.34}}$ $\mathbf{79.6^{\pm 0.37}}$ $\mathbf{83.9^{\pm 0.36}}$ $84.4^{\pm 0.15}$ $\mathbf{66.5^{\pm 0.24}}$ $89.7^{\pm 0.07}$ $\mathbf{68.4^{\pm 0.38}}$ $64.1^{\pm 0.09}$ $\mathbf{72.7^{\pm 0.17}}$
Last
   InfLoRA $18.26^{\pm 0.49}$ $\mathbf{82.36^{\pm 0.92}}$ $46.57^{\pm 0.89}$ $79.38^{\pm 2.22}$ $76.16^{\pm 1.61}$ $79.58^{\pm 0.60}$ $95.74^{\pm 0.44}$ $87.78^{\pm 0.85}$ $71.11^{\pm 0.73}$ $73.05^{\pm 0.19}$ $70.99^{\pm 0.24}$
   CoDyRA $\mathbf{37.7^{\pm 0.42}}$ $81.5^{\pm 0.24}$ $65.1^{\pm 0.63}$ $89.9^{\pm 0.55}$ $91.4^{\pm 0.38}$ $85.5^{\pm 0.16}$ $96.8^{\pm 0.08}$ $\mathbf{93.3^{\pm 0.30}}$ $77.3^{\pm 0.66}$ $73.5^{\pm 0.21}$ $79.2^{\pm 0.18}$
   MoRA $\mathbf{37.7^{\pm 0.28}}$ $81.5^{\pm 0.22}$ $\mathbf{70.7^{\pm 0.49}}$ $\mathbf{92.4^{\pm 0.20}}$ $\mathbf{95^{\pm 0.34}}$ $\mathbf{86^{\pm 0.13}}$ $\mathbf{97.6^{\pm 0.19}}$ $92.6^{\pm 0.10}$ $\mathbf{81^{\pm 0.35}}$ $\mathbf{74.7^{\pm 0.06}}$ $\mathbf{80.9^{\pm 0.12}}$
A.6 More comparisons on X-TAIL
To further validate MoRA’s robustness, we evaluated it under other continual-learning task orderings, i.e., X-TAIL (Order 2), as shown in Table 14. The results align with those in Table 1, confirming that MoRA consistently achieves state-of-the-art performance.
Table 14: Comparisons on X-TAIL (Order 2) for each domain in terms of “Transfer”, “Average”, and “Last” scores (%).
Method
Cars Aircraft OxfordPet Food SUN397 MNIST Flowers DTD Caltech101 EuroSAT Average
CLIP
   Zero-shot 66.1 23.5 86.7 84 63.7 46.7 63.6 37.3 76.8 36.7 58.5
Transfer
   Zero-shot (Radford et al., 2021) 23.5 86.7 84 63.7 46.7 63.6 37.3 76.8 36.7 57.7
   LwF (Li & Hoiem, 2017) 20.0 74.1 79.6 58.1 34.1 48.9 27.7 64.4 15.1 46.9
   WiSE-FT (Wortsman et al., 2022) 21.3 79.5 83.3 61.0 39.9 56.5 29.6 68.0 20.8 51.1
   ZSCL (Zheng et al., 2023) 23.0 84.3 87.2 63.0 42.1 65.2 34.6 71.4 40.9 56.9
   MoE-Adapter (Yu et al., 2024) 17.1 87.2 87.5 58.4 12.6 65.5 35.9 70.0 17.9 50.2
   RAIL-Primal (Xu et al., 2024) 23.5 86.7 84 63.7 46.7 63.6 37.3 76.8 36.7 57.7
   CoDyRA (Lu et al., 2024) 23.6 89.2 83 62 51 71.4 38 77.4 39 59.4
   MoRA 23.6 88.7 83.4 62.6 51.2 69.9 39.3 77.5 39 59.5
Average
   LwF (Li & Hoiem, 2017) 49.0 27.4 69.7 83.0 65.7 42.2 63.5 33.1 68.5 17.5 52.0
   WiSE-FT (Wortsman et al., 2022) 57.9 29.6 77.8 85.4 68.0 51.6 69.3 35.5 71.0 23.0 56.9
   ZSCL (Zheng et al., 2023) 74.4 36.4 86.7 88.7 68.9 50.0 75.1 40.1 72.5 43.7 63.6
   MoE-Adapter (Yu et al., 2024) 74.4 38.6 87.7 87.3 67.9 50.6 76.5 43.7 72.3 18.8 61.8
   RAIL-Primal (Xu et al., 2024) 77.9 40.4 85.6 83.3 68.3 62.2 76.6 45.8 80.4 41.7 66.2
   CoDyRA (Lu et al., 2024) 80 39.2 92.5 85.2 69.2 73.7 79.6 46.2 78.6 44.1 68.8
   MoRA 80.2 40.1 92.5 84.7 70.1 74 80.1 48.7 78.4 44.4 69.3
Last
   LwF (Li & Hoiem, 2017) 29.6 17.5 63.0 83.8 67.7 44.9 79.3 44.8 84.6 39.0 55.4
   WiSE-FT (Wortsman et al., 2022) 46.1 23.5 71.3 85.7 70.2 59.1 85.5 47.9 82.4 42.8 61.5
   ZSCL (Zheng et al., 2023) 71.7 35.3 86.5 89.2 71.8 52.3 89.8 52.0 77.1 68.4 69.4
   MoE-Adapter (Yu et al., 2024) 75.1 41.1 87.9 87.1 74.1 89.7 92.6 61.2 81.0 27.4 71.7
   RAIL-Primal (Xu et al., 2024) 77.7 41.9 86.1 83.3 71.8 91.6 97.3 66.4 94.8 86.9 79.8
   CoDyRA (Lu et al., 2024) 79 38.6 92.6 86.4 74.7 95.2 93 64.7 81.9 92.2 79.8
   MoRA 79.3 38.9 93.1 85.4 74.9 96.4 94.1 69.9 82 92.9 80.7
Table 15: Trainable parameters and averaged training GPU memory per task.
Method Trainable Params. (Million) GPU Mem. (MiB)
LWF (Li & Hoiem, 2017) 129.6 32172
ZSCL (Zheng et al., 2023) 129.6 26290
MoE-Adapters (Yu et al., 2024) 59.8 22358
CoDyRA (Lu et al., 2024) 4.4 21770
MoRA 4.4 21090
A.7 Computation Cost